Someday a person will walk on Mars, which is 5.6 x 1010 m from the earth at the closest approach. Determine the minimum time that will be required for a message from Mars to reach the earth via radio waves.


1
Expert's answer
2021-11-09T17:53:14-0500

"t=\frac{2s}c,"

"t=\frac{2\cdot 5.6\cdot 10^{10}}{3\cdot 10^8}=18.7^m."